OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (studygurl14):

speed = distance/time D = distance, S = speed, T = time, A = car A, B = car B \[D_A=D_B\] \[S_A=\frac{ D_A }{ T_A } \rightarrow S_A=\frac{ D_A }{ 2.4 }\] \[S_A = S_B+22\] Can you do the rest?

OpenStudy (anonymous):

@StudyGurl14 is it 2.4+22?

OpenStudy (studygurl14):

hold on a sec

OpenStudy (anonymous):

k :)

OpenStudy (anonymous):

@StudyGurl14 ?

OpenStudy (studygurl14):

hold on. everyone keeps linking my name and asking me thinks. did you try googling it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

someone said 33mph

OpenStudy (studygurl14):

I missed something. \[S_B=\frac{ D_B }{ 4 }\] That should help. try again

OpenStudy (studygurl14):

i think the answer is 55 though, not sure....

OpenStudy (studygurl14):

Yep the answer is 55mph http://openstudy.com/study#/updates/542194d4e4b0e08980376271